OpenStudy (anonymous):

last one-- The length of a rectangle is 3 feet less than twice the width. The perimeter is 84 feet. Find the length and width.

OpenStudy (anonymous):

length l, width w l= 2w-3 perimeter 2(l+w) = 84 2(2w-3+w)=84 -6+6w=84 w=90/6= 15 width=15, length= 27

OpenStudy (anonymous):

thanks born2die20!